OpenStudy (thewhovian33):

What is the square root of -80? Options: -4√5 -4√ 5i 4√ 5i 4√ 5

TheSmartOne (thesmartone):

Hint: \(\sqrt{-1} = i\)

MARC (marc_d):

learnt new stuff.. :)

OpenStudy (luffingsails):

Think of \[\sqrt{80}=\sqrt{-1*80}=\sqrt{-1}*\sqrt{80}\]

OpenStudy (luffingsails):

Then use @TheSmartOne's note to solve.

OpenStudy (acespeedfighter):

🔴҈҈҈҈̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚҉̚